# Multi-Atlas Brain Network Classification through Consistency Distillation and Complementary Information Fusion

### v1

@@ Abstract
In the realm of neuroscience, identifying distinctive patterns associated with neurological disorders via brain networks is crucial. Resting-state functional magnetic resonance imaging (fMRI) serves as a primary tool for mapping these networks by correlating blood-oxygen-level-dependent (BOLD) signals across different brain regions, defined as regions of interest (ROIs). Constructing these brain networks involves using atlases to parcellate the brain into ROIs based on various hypotheses of brain division. However, there is no standard atlas for brain network classification, leading to limitations in detecting abnormalities in disorders. Some recent methods have proposed utilizing multiple atlases, but they neglect consistency across atlases and lack ROI-level information exchange. To tackle these limitations, we propose an Atlas-Integrated Distillation and Fusion network (AIDFusion) to improve brain network classification using fMRI data. AIDFusion addresses the challenge of utilizing multiple atlases by employing a disentangle Transformer to filter out inconsistent atlas-specific information and distill distinguishable connections across atlases. It also incorporates subject- and population-level consistency constraints to enhance cross-atlas consistency. Additionally, AIDFusion employs an inter-atlas message-passing mechanism to fuse complementary information across brain regions. Experimental results on four datasets of different diseases demonstrate the effectiveness and efficiency of AIDFusion compared to state-of-the-art methods. A case study illustrates AIDFusion extract patterns that are both interpretable and consistent with established neuroscience findings.
1 Introduction
In the field of neuroscience, a key objective is to identify distinctive patterns associated with neurological disorders (e.g., Alzheimer’s, Parkinson’s, and Autism) by the brain networks [42]. Resting-state functional magnetic resonance imaging (fMRI) is widely employed among various neuroimaging techniques to characterize the connectivities among brain regions [53]. This results in brain networks where each node represents a specific brain region, referred to as a region of interest (ROI). Each edge indicates a pairwise correlation between the blood-oxygen-level-dependent (BOLD) signals of two ROIs [54], revealing the connectivity between brain regions and indicating which areas tend to be activated synchronously or exhibit correlated activities.
@@ 1 Introduction
  • We propose a multi-atlas solution for brain network classification with fMRI data. AIDFusion takes full advantage of multi-atlas brain networks by enhanced atlas-consistent information distillation and intense fusion of cross-atlas complementary information.
  • We evaluate AIDFusion on four resting-state fMRI brain network datasets for different neurological disorders. Our results demonstrate the superiority of AIDFusion over state-of-the-art baseline methods in terms of effectiveness and efficiency in brain network classification.
  • We present a case study that underscores the intriguing, straightforward, and highly interpretable patterns extracted by our approach, aligning with domain knowledge found in neuroscience literature.
2 Related Work
2.1 Brain Network Analysis with Various Atlases
@@
Multi-atlas methods introduce multiple brain atlases for each neuroimage, which can provide information that complements each other and offers ample details without being restricted by the parcellation mode. MGRL [7] pioneered the construction of multi-atlas brain networks using various atlases. It applied graph convolutional networks (GCNs) to learn multi-atlas representations and perform graph-level fusion for disease classification. METAFormer [35] proposed a multi-atlas enhanced transformer approach with self-supervised pre-training for Autism spectrum disorder (ASD) classification. A graph-level late fusion was utilized to aggregate the representations of different atlases. Lee et al. [27] employed a multi-atlas fusion approach that integrates early fusion on the raw feature to capture complex brain network patterns. STW-MHGCN [29] constructs a spatial and temporal weighted hyper-connectivity network to fuse multi-atlas information, and Huang et al. [17] adopt a voting strategy to integrate the classification results of different classifiers (each corresponding to a different atlas) for ASD diagnosis. However, these studies did not consider the inherent consistency between atlases. Independently encoding multi-atlas brain networks without constraints might extract atlas-specific information, distracting from disease-related pattern modeling. Moreover, existing works only incorporate primitive early or late feature fusion between atlases. This absence of intermediate ROI-level interaction could hinder their models’ ability to discern complementary information in each atlas. To the best of our knowledge, our work is the first to introduce information distillation with consistency constraints and employ intermediate ROI-level interaction for complementary information fusion. Note that in our work, multiple atlases are applied to preprocessed images for parcellation, meaning our method is based on a single template. The difference between multi-atlas methods and multi-template methods are discussed in Appendix A.
@@
Multi-modal and multi-resolution methods also explore brain networks using various atlases. Research about multi-modal brain networks [63, 64, 26, 57, 65] employed multiple modalities of neuroimaging data, including fMRI, Diffusion Tensor Imaging (DTI) and Positron Emission Tomography (PET), with various atlases to enhance brain network classification, as different modalities provide abundant information compared to a single modality. However, these multi-modal methods focus on fusing structural and functional connectivity information instead of trying to capture the whole picture of the single modality data. Another line of research [32, 33, 52] focuses on applying multi-resolution atlases to fMRI data to capture individual behavior across coarse-to-fine scales. However, the technical design of these approaches focuses on extracting information from both fine and coarse scales under the same parcellation mode. Although multi-modal and multi-resolution methods employ various atlases, they focus on different objectives from multi-atlas approaches, and the field of brain network analysis with multi-atlas is still in its infancy stage.
2.2 Graph Neural Networks (GNNs)
@@
In recent years, there has been a surge of interest in employing Graph Neural Networks (GNNs) for the analysis of brain networks. Ktena et al. [24] utilized graph convolutional networks to learn similarities between pairs of brain networks (subjects). BrainNetCNN [20] introduced edge-to-edge, edge-to-node, and node-to-graph convolutional filters to harness the topological information within brain networks. MG2G [56] utilizes a two-stage method where it initially learns node representations using an unsupervised stochastic graph embedding model based on latent distributions, which are then used to train a classifier, allowing for the identification of significant ROIs with Alzheimer’s disease (AD)-related effects. Zhang et al. [62] incorporated both local ROI-GNN and global subject-GNN guided by non-imaging data. ContrastPool [55] introduced a dual attention mechanism to extract discriminative features across ROIs for subjects within the same group.
@@
An alternative method for graph representation learning involves Transformer-based models [48], which adapt the attention mechanism to consider global information for each node and incorporate positional encoding to capture graph topological information. Graph Transformers have garnered significant attention due to their impressive performance in graph representation learning [15, 59, 43]. A series of brain network Transformer methods have emerged for brain network analysis. One such method [18] applied Transformers to learn pairwise connection strengths among brain regions across individuals. THC [12] introduced an interpretable Transformer-based model for joint hierarchical cluster identification and brain network classification. DART [19] utilized segmenting BOLD signals to generate dynamic brain networks and then incorporated them with static networks for representation learning. Most GNN- and Transformer-based methods for brain network analysis are designed for single-atlas thus may lead to a dependency on specific parcellation mode.
3 Preliminaries
3.1 Brain Network Construction
@@
3.2 Problem Definition
Multi-atlas brain network classification aims to predict the distinct class of each subject by using various atlases for the same fMRI data. Given a dataset of labeled brain networks $\mathcal{D}=\{(\boldsymbol{X}^{a},\boldsymbol{X}^{b},y_{X})\}$, where $y_{X}$ is the class label of brain networks $\boldsymbol{X}^{a}$ and $\boldsymbol{X}^{b}$, the problem of brain network classification is to learn a predictive function $f$: $(\boldsymbol{X}^{a},\boldsymbol{X}^{b})\rightarrow y_{X}$, which maps input brain networks to the groups they belong to, expecting that $f$ also works well on unseen brain networks. We summarize the notations used throughout the paper in Appendix B.
Figure 1: The framework of AIDFusion for multi-atlas brain network classification.
4 Methodology
In this section, we provide a detailed exposition of the design of our proposed Atlas-Integrated Distillation and Fusion network (AIDFusion), depicted in Figure 1. Two brain networks constructed with different atlases are separately processed in our model. In the following, we first introduce the disentangle Transformer with identity embedding to remove inconsistent atlas-specific information (Section 4.1). We then describe the inter-atlas message-passing for spatial-based intense fusion of cross-atlas (Section 4.2). Finally, we discuss our design of the losses that enforce atlas-consistent information distillation with domain considerations (Section 4.3).
4.1 Disentangle Transformer with Identity Embedding
Identity Embedding. In graph Transformer models, positional embedding is commonly used to encode the topological information of the graph. However, designs like distance-based, centrality-based, and eigenvector-based positional embeddings [28, 59, 49] are impractical for brain networks due to their high density (always fully connected). Correlation-based brain networks already contain sufficient positional information for ROIs, making general positional embeddings both costly and redundant. Instead, we propose a learnable identity embedding that adaptively learns a unique identity for each ROI, aligning nodes in the same ROI across the same atlas. This embedding assigns the same identity to nodes within the same ROI. As shown in Eq. (1), we introduce a parameter matrix $\boldsymbol{W}_{ID}$ to encode node identities alongside original node features $\boldsymbol{X}$, with $\operatorname{MLP}(\cdot)$ denoting a multilayer perceptron (MLP).
@@ 4.1 Disentangle Transformer with Identity Embedding
[…]
@@
$\small\mathcal{L}_{orth}=\frac{1}{r}\sum\frac{||\boldsymbol{W}_{INC}^{a}\cdot\boldsymbol{W}_{INC}^{b}||}{||\boldsymbol{W}_{INC}^{a}||\cdot||\boldsymbol{W}_{INC}^{b}||}.$ (5)
4.2 Inter-Atlas Message-Passing
The features at different atlases originate from totally different parcellation modes. Pulling those highly correlated features of two different atlases into a shared space allows their effective fusion. Existing literature on multi-atlas brain networks independently learns the representations of ROIs in each atlas without exchanging information across atlases [7, 27]. Additionally, the spatial relationship between ROIs in different atlases is neglected in these works. Our proposed AIDFusion enables inter-atlas message-passing between neighboring regions in different atlases by considering spatial information. Specifically, we use the spatial distance between the centroids of ROIs in different atlases to construct inter-atlas connections. As shown in Figure 1c, we utilize the $k$-nearest-neighbor ($k$NN) algorithm to connect each ROI to $k$ ROIs from the other atlas. Note that we only construct inter-atlas connections without considering intra-atlas connections since the information exchange within the same atlas has already proceeded in previous disentangle Transformer. Afterwards, an adjacency matrix $\boldsymbol{A}^{ab}\in\{0,1\}^{(n^{a}+n^{b})\times(n^{a}+n^{b})}$ is obtained and used for graph convolution [22]:
@@ 4.2 Inter-Atlas Message-Passing
where $\sigma$ is the activation function (e.g., ReLU), $\boldsymbol{D}$ is the degree matrix of $\boldsymbol{A}^{ab}$, $\boldsymbol{H}^{ab}\in\mathbb{R}^{(n^{a}+n^{b})\times d}$ is the combined node representation matrix for the two atlases, and $\boldsymbol{W}_{GC}$ is the learnable weight matrix of the GCN layer. An example of the adjacency matrix $\boldsymbol{A}^{ab}$ that used for inter-atlas message-passing is shown and discussed in Appendix C.
4.3 Subject- and Population-level Consistency
Subject-level Consistency. To ensure the high-level consistency for the two brain networks from different atlases, we introduce a contrastive loss on the subject level. First, we apply DiffPool [60] to each atlas to capture higher-level patterns. The DiffPool contains two GCN layers. $\operatorname{GCN}_{pool}$ is used to learn a cluster assignment matrix $\boldsymbol{S}\in\mathbb{R}^{n\times n^{\prime}}$ as shown in Eq. (7). Herein, $n^{\prime}$ is a pre-defined number of clusters controlled by a hyperparameter named the pooling ratio. The other $\operatorname{GCN}_{emb}$ is used to obtain the embedded node feature matrix $\boldsymbol{Z}\in\mathbb{R}^{n\times d}$ as shown in Eq. (8). Both these two GCNs are defined similarly with Eq. (6). We sparsify the connectivity matrices $\boldsymbol{X}$ by keeping top 20% correlations and use them as the adjacency matrices $\boldsymbol{A}$ in these two GCNs, to avoid over-smoothing. The feature matrices of two atlas $\boldsymbol{H}$ are obtained from the output of inter-atlas message-passing.

### v2

@@ Abstract
Brain network analysis plays a crucial role in identifying distinctive patterns associated with neurological disorders. Functional magnetic resonance imaging (fMRI) enables the construction of brain networks by analyzing correlations in blood-oxygen-level-dependent (BOLD) signals across different brain regions, known as regions of interest (ROIs). These networks are typically constructed using atlases that parcellate the brain based on various hypotheses of functional and anatomical divisions. However, there is no standard atlas for brain network classification, leading to limitations in detecting abnormalities in disorders. Recent methods leveraging multiple atlases fail to ensure consistency across atlases and lack effective ROI-level information exchange, limiting their efficacy. To address these challenges, we propose the Atlas-Integrated Distillation and Fusion network (AIDFusion), a novel framework designed to enhance brain network classification using fMRI data. AIDFusion introduces a disentangle Transformer to filter out inconsistent atlas-specific information and distill meaningful cross-atlas connections. Additionally, it enforces subject- and population-level consistency constraints to improve cross-atlas coherence. To further enhance feature integration, AIDFusion incorporates an inter-atlas message-passing mechanism that facilitates the fusion of complementary information across brain regions. We evaluate AIDFusion on four resting-state fMRI datasets encompassing different neurological disorders. Experimental results demonstrate its superior classification performance and computational efficiency compared to state-of-the-art methods. Furthermore, a case study highlights AIDFusion’s ability to extract interpretable patterns that align with established neuroscience findings, reinforcing its potential as a robust tool for multi-atlas brain network analysis. The code is publicly available at https://github.com/AngusMonroe/AIDFusion.
I Introduction
In the field of neuroscience, a key objective is to identify distinctive patterns associated with neurological disorders (e.g., Alzheimer’s, Parkinson’s, and Autism) by the brain networks [1, 2]. Resting-state functional magnetic resonance imaging (fMRI) is widely employed among various neuroimaging techniques to characterize the connectivities among brain regions [3]. This results in brain networks where each node represents a specific brain region, referred to as a region of interest (ROI). Each edge indicates a pairwise correlation between the blood-oxygen-level-dependent (BOLD) signals of two ROIs [4], revealing the connectivity between brain regions and indicating which areas tend to be activated synchronously or exhibit correlated activities.
@@ I Introduction
Unlike existing multi-atlas methods that typically apply early or late fusion heuristics, AIDFusion introduces an end-to-end architecture capable of learning consistent and complementary representations from heterogeneous atlases. It achieves this with fewer parameters, faster convergence, and better generalization performance, particularly in out-of-distribution (OOD) scenarios such as cross-site evaluations.
  • A novel architecture for multi-atlas brain network classification. AIDFusion integrates disentangled representation learning, inter-atlas message passing, and multi-level consistency constraints to effectively leverage complementary information from multiple brain atlases.
  • Comprehensive evaluation across four resting-state fMRI datasets. AIDFusion consistently outperforms 11 strong baselines, including both classical ML, general-purposed GNN and brain network modes, not only in accuracy but also in training efficiency.
  • Interpretability and robustness. AIDFusion extracts highly interpretable connectivity patterns that align with known neuroscience findings and demonstrates strong generalization in cross-site experiments, highlighting its clinical applicability and robustness to domain shifts.
II Related Work
II-A Brain Network Analysis with Various Atlases
@@
Multi-atlas methods introduce multiple brain atlases for each neuroimage, which can provide information that complements each other and offers ample details without being restricted by the parcellation mode. MGRL [12] pioneered the construction of multi-atlas brain networks using various atlases. It applied graph convolutional networks (GCNs) to learn multi-atlas representations and perform graph-level fusion for disease classification. METAFormer [13] proposed a multi-atlas enhanced transformer approach with self-supervised pre-training for Autism spectrum disorder (ASD) classification. A graph-level late fusion was utilized to aggregate the representations of different atlases. Lee et al. [14] employed a multi-atlas fusion approach that integrates early fusion on the raw feature to capture complex brain network patterns. STW-MHGCN [15] constructs a spatial and temporal weighted hyper-connectivity network to fuse multi-atlas information, and Huang et al. [16] adopt a voting strategy to integrate the classification results of different classifiers (each corresponding to a different atlas) for ASD diagnosis. CcSi-MHAHGEL [17] introduces a class-consistency and site-independence Multiview Hyperedge-Aware HyperGraph Embedding Learning framework to integrate brain networks constructed on multiple atlases in a multisite fMRI study. However, these studies did not consider the inherent consistency between atlases. Independently encoding multi-atlas brain networks without constraints might extract atlas-specific information, distracting from disease-related pattern modeling. Moreover, existing works only incorporate primitive early or late feature fusion between atlases. This absence of intermediate ROI-level interaction could hinder their models’ ability to discern complementary information in each atlas. To the best of our knowledge, our work is the first to introduce information distillation with consistency constraints and employ intermediate ROI-level interaction for complementary information fusion. Note that in our work, multiple atlases are applied to preprocessed images for parcellation, meaning our method is based on a single template.
@@
Multi-modal and multi-resolution methods also explore brain networks using various atlases. Research about multi-modal brain networks [18, 19, 20, 21, 22, 23] employed multiple modalities of neuroimaging data, including fMRI, Diffusion Tensor Imaging (DTI) and Positron Emission Tomography (PET), with various atlases to enhance brain network classification, as different modalities provide abundant information compared to a single modality. However, these multi-modal methods focus on fusing structural and functional connectivity information instead of trying to capture the whole picture of the single modality data. Another line of research [24, 25, 26, 27] focuses on applying multi-resolution atlases to fMRI data to capture individual behavior across coarse-to-fine scales. However, the technical design of these approaches focuses on extracting information from both fine and coarse scales under the same parcellation mode. Although multi-modal and multi-resolution methods employ various atlases, they focus on different objectives from multi-atlas approaches, and the field of brain network analysis with multi-atlas is still in its infancy stage.
II-B Graph Neural Networks (GNNs)
@@
In recent years, Graph Neural Networks (GNNs) have gained significant attention for brain network analysis. Ktena et al. [28] applied graph convolutional networks to learn similarities between pairs of brain networks (subjects). BrainNetCNN [5] introduced edge-to-edge, edge-to-node, and node-to-graph convolutional filters to capture the spatial information within brain networks. MG2G [29] employs a two-stage approach, where node representations are initially learned using an unsupervised stochastic graph embedding model based on latent distributions. These representations are then used to train a classifier for identifying significant ROIs associated with Alzheimer’s disease (AD). Zhang et al. [30] combined local ROI-GNN and global subject-GNN models to learn the brain network representations. ContrastPool [31] introduced a dual-attention mechanism to extract discriminative features across ROIs for subjects within the same group.
@@
An alternative approach for graph representation learning involves Transformer-based models [32], which adapt the attention mechanism to incorporate global information for each node and positional encoding to capture graph topology. Graph Transformers have gained significant attention due to their impressive performance in graph representation learning [33, 34, 35]. Several Transformer-based methods have emerged for brain network analysis. For example, one method [36] applied Transformers to learn pairwise connection strengths among brain regions across individuals. Com-BrainTF [37] uses a hierarchical local-global transformer to generate community-aware node embeddings, while GBT [38] employs an attention weight matrix approximation to focus on the most relevant components for enhanced graph representation. THC [39] introduced an interpretable Transformer-based model for joint hierarchical cluster identification and brain network classification. DART [40] segmented BOLD signals to generate dynamic brain networks, which were then integrated with static networks for improved representation learning. Contrasformer [41] introduces a contrast graph to highlight the difference between groups and thus improve the model’s generalization ability. Most GNN- and Transformer-based methods for brain network analysis are designed for single-atlas use, which may create a dependency on specific parcellation schemes.
Figure 1: The framework of AIDFusion for multi-atlas brain network classification. The proposed framework includes three key components: Disentangle Transformer, Inter-Atlas Message-Passing, and Subject- and Population-level Consistency Constraint.
III Preliminaries
III-A Brain Network Construction
@@
III-B Problem Definition
Multi-atlas brain network classification aims to predict the distinct class of each subject by using brain networks constructed with various atlases for the same fMRI data. Given a labeled brain network dataset $\mathcal{D}=\{(\boldsymbol{X}^{a},\boldsymbol{X}^{b},y_{X})\}$, where $y_{X}$ is the class label of brain networks $\boldsymbol{X}^{a}$ and $\boldsymbol{X}^{b}$, the objective of multi-atlas brain network classification is to learn a predictive function $f$: $(\boldsymbol{X}^{a},\boldsymbol{X}^{b})\rightarrow y_{X}$ that can map the input brain networks to the groups they belong to, expecting that the function $f$ also works well on the other unseen brain networks.
IV Methodology
In this section, we provide a detailed exposition of the design of our proposed Atlas-Integrated Distillation and Fusion network (AIDFusion), depicted in Figure 1. Two brain networks constructed with different atlases are separately processed in our model. In the following, we first introduce the disentangle Transformer with identity embedding to remove inconsistent atlas-specific information (Section IV-A). We then describe the inter-atlas message-passing for spatial-based intense fusion of cross-atlas (Section IV-B). Finally, we discuss our design of the losses that enforce atlas-consistent information distillation with domain considerations (Section IV-C).
IV-A Disentangle Transformer with Identity Embedding
Identity Embedding. In graph Transformer models, positional embedding is commonly used to encode the topological information of the graph. However, designs like distance-based, centrality-based, and eigenvector-based positional embeddings [43, 34, 44] are impractical for brain networks due to their high density (always fully connected). Correlation-based brain networks already contain sufficient positional information for ROIs, making general positional embeddings both costly and redundant. Instead, we propose a learnable identity embedding that adaptively learns a unique identity for each ROI, aligning nodes in different subjects that correspond to the same ROI in the same atlas. This embedding assigns the same identity to nodes within the same ROI. The design is inspired by the concept of word embeddings in natural language processing [45], where each word is assigned an independent learnable representation. As shown in Eq. (1), we introduce a parameter matrix $\boldsymbol{W}_{ID}$ to encode node identities alongside original node features $\boldsymbol{X}$, with $\operatorname{MLP}(\cdot)$ denoting a multilayer perceptron (MLP).
@@ IV-A Disentangle Transformer with Identity Embedding
where $\left[\begin{array}[]{c}\cdot\\ \cdot\\ \end{array}\right]$ denotes the append operation. To enforce each incompatible node captures different information, we initialize them using the Gram-Schmidt process [48] to ensure they are orthogonal to each other. Then the self-attention function [32] is applied to $\boldsymbol{H}^{\prime}\in\mathbb{R}^{(n+r)\times d}$:
@@
[…]
IV-B Inter-Atlas Message-Passing
The features in different atlases originate from totally different parcellation modes. Pulling those highly correlated features of two different atlases into a shared space allows their effective fusion. Existing literature on multi-atlas brain networks independently learns the representations of ROIs in each atlas without exchanging information across atlases [12, 14]. Additionally, the spatial relationship between ROIs in different atlases is neglected in these works. Our proposed AIDFusion enables inter-atlas message-passing between neighboring regions in different atlases by considering spatial information. Specifically, for each ROI, we define its spatial coordinate as the centroid of the voxels assigned to that region and compute the Euclidean distance between ROIs in different atlases as the spatial distance. As shown in Figure 1c, we utilize the $k$-nearest-neighbor (kkNN) algorithm to connect each ROI to $k$ ROIs from the other atlas. Note that we only construct inter-atlas connections without considering intra-atlas connections since the information exchange within the same atlas has already been processed in the previous disentangle Transformer. When atlases are derived from different templates (e.g., MNI vs. Talairach), coordinate alignment becomes a critical issue. Fortunately, spatial normalization and template conversion between standard spaces is a well-established task in neuroimaging. Several mature techniques [49, 50, 51, 52] exist to handle such transformations accurately. These methods can be readily incorporated into our framework as a preprocessing step to ensure spatial compatibility before applying inter-atlas message-passing.
Afterwards, an adjacency matrix $\boldsymbol{A}^{ab}\in\{0,1\}^{(n^{a}+n^{b})\times(n^{a}+n^{b})}$ is obtained and used for graph convolution [53]:
@@ IV-B Inter-Atlas Message-Passing
where $\sigma$ is the activation function (e.g., ReLU), $\boldsymbol{D}$ is the degree matrix of $\boldsymbol{A}^{ab}$, $\boldsymbol{H}^{ab}\in\mathbb{R}^{(n^{a}+n^{b})\times d}$ is the combined node representation matrix for the two atlases, and $\boldsymbol{W}_{GC}$ is the learnable weight matrix of the GCN layer.
IV-C Subject- and Population-level Consistency
Subject-level Consistency. To ensure the high-level consistency for the two brain networks from different atlases, we introduce a contrastive loss on the subject level. First, we apply DiffPool [54] to each atlas to capture higher-level patterns. The DiffPool contains two GCN layers. $\operatorname{GCN}_{pool}$ is used to learn a cluster assignment matrix $\boldsymbol{S}\in\mathbb{R}^{n\times n^{\prime}}$ as shown in Eq. (7). Herein, $n^{\prime}$ denotes the number of clusters, which is controlled by a preset pooling ratio. The other $\operatorname{GCN}_{emb}$ is used to generate the embedded feature $\boldsymbol{Z}\in\mathbb{R}^{n\times d}$ as shown in Eq. (8). Both these two GCNs are defined similarly with Eq. (6). We sparsify the connectivity matrices $\boldsymbol{X}$ by keeping top 20% correlations and use them as the adjacency matrices $\boldsymbol{A}$ in these two GCNs, to avoid over-smoothing. The feature matrices of two atlas $\boldsymbol{H}$ are obtained from the output of inter-atlas message-passing.